k

Solve the differential equations using the method of Variation of Parameters: 2y - y - y=tet UTICA
Answer #1

zy - y-y = te : 2 mm-120 > (2m +1)(m-1) = 0 am= -1/2 1 Ye=cé teet - Y=et w*.*.-en set set .: H+) = tek Y2GH W(4,4) dt + et1/4 tet 12a e 49(6) W($1, I2) dt dt = 2 2 St at =름(뷔) = 북 ..Lp = y co, te e t z et 46 9 e + face ..Y=ae t } + C₂ e